OpenStudy (anonymous):

Solve: -(3x - 2) = -4(x + 1) - 4

OpenStudy (anonymous):

simplify or answer?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

-(3x-2)=-4(x+1)-4 -3x+2=-4x-4-4 -3x+4x+2+4+4=0 x+10=0 x=-10 Check: substitute x=-10 into equation -(3x-2)=-3(-10)+2=32 -4(x+1)-4 = -4(-10+1)-4=-4(-9)-4=36-4=32 Both sides equal so x=-10 is correct

OpenStudy (anonymous):

Thankyou @cutepochacco

OpenStudy (anonymous):

your welcome